Arna is twice as old as Mae. In 4 years time, she will be five-thirds as old as Mae. How old is each now

with explenation pls
1
Expert's answer
2019-09-20T11:15:49-0400

Let a be the age of Arna, m be the age of Mae.

"a=2m, \;\;a+4=\frac{5}{3}(m+4)"

Then, "2m+4=\frac{5}{3}(m+4)"

"\frac{1}{3}m=\frac{8}{3}"

m=8, a=2m=16.

Arna is 16 years old, Mae is 8 years old.